OpenStudy (nathan917):

A bag contains 3 blue marbles, 9 green marbles, and 11 yellow marbles. Twice you draw a marble and replace it. Find P(blue, then green). twenty-seven over five hundred twenty-nine twenty-seven over twenty-three fifteen over five hundred twenty-nine twelve over twenty-three

OpenStudy (nathan917):

(3/23) * (9/23) = 27/529?

OpenStudy (nathan917):

@tcarroll010

OpenStudy (rulnick):

P(blue then green) = P(blue) P(green) = (3/23) (9/23) = ... so yes

OpenStudy (nathan917):

Thanks.

OpenStudy (rulnick):

welcome